5. What is the real future value of $500 after 10 years if the nominal interest rate is 13% and inflation is 2%? (5)
Answer #1
Real interest rate =((1 + Nominal interest rate) / (1+ Inflation rate))-1
Real interest rate = ((1+13%)/(1+2%))-1 = 10.78%
FV = PV * (1+interest rate)^number of years
Real Future value = 500*((1+10.78%)^10) = $ 1,391.82
Real Future value = $ 1,391.82